# Properties of observable mixed inertial/gravito-inertial modes in $\gamma$ Doradus stars

### v1

@@ Abstract
Context. The space missions Kepler and TESS provided a large number of highly detailed time series for main sequence stars, including $\gamma$ Doradus ($\gamma$ Dor) stars. Additionally, numerous $\gamma$ Dor stars are to be observed in the near future thanks to the upcoming PLATO mission. In $\gamma$ Doradus stars, gravito-inertial modes in the radiative zone and inertial modes in the convective core can interact resonantly, which translates into the appearance of dip structures in the period spacing of modes. Those dips are information-rich as they are related to the star core characteristics.
Aims. We aim at characterising these dips according to stellar properties and thus developing new seismic diagnostic tools to constrain the internal structure of $\gamma$ Dor stars, especially their core.
Methods. We used the two-dimensional oscillation code TOP to compute sectoral prograde and axisymmetric dipolar modes in $\gamma$ Dor stars at different rotation rates and evolutionary stages. We then characterised the dips we obtained by their width and location on the period spacing diagram.
Results. We found that the width and the location of the dips depend quasi-linearly on the ratio of the rotation rate and the Brunt-Väisälä frequency at the core interface. This allowed us to determine empirical relations between width and location of dips on the one hand, and on the other, resonant inertial mode frequency in the core and Brunt-Väisälä frequency at the core/radiative zone interface. We also propose an approximate theoretical model to support and discuss these empirical relations.
Conclusions. The empirical relations we established could be applied to dips observed in data, which would allow for the estimate of frequencies of resonant inertial modes in the core and of the Brunt-Väisälä jump at the interface between the core and the radiative zone. As those two parameters are both related to the evolutionary stage of the star, their determination could lead to more accurate estimations of stellar ages.
@@ 1 Introduction
Asteroseismology has proven to be a crucial tool in unravelling the mysteries of stellar evolution and determining the ages and evolutionary stages of stars (e.g. Lebreton et al., 2014; Aerts, 2021, and references therein). Estimating the evolutionary stages of stars is particularly significant in addressing the complex question of angular momentum transport within stars (Aerts et al., 2019).
@@
Among pulsating stars, $\gamma$ Doradus (hereafter $\gamma$ Dor) stars are main sequence stars that have typically a mass of 1.4–2.0 M and a radius of 1.4–2.3 R (Kaye et al., 1999; Uytterhoeven et al., 2011). Their structure is characterised by a convective core, a large radiative zone and a small sub-surface convective envelope. They exhibit non-radial gravity modes oscillating with a period of 0.3 to 3 d (Kaye et al., 1999). Thanks to high-precision photometry provided by space-borne instruments, especially by the Kepler mission (Borucki et al., 2010) and the Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2014), a large quantity of good quality data is now available and enables for the study of numerous $\gamma$ Dor stars.
@@
Gravito-inertial modes have been identified in 611 $\gamma$ Dor stars observed by Kepler (Li et al., 2020) and in 106 $\gamma$ Dor stars observed by TESS (Garcia et al., 2022). They enabled to determine both the rotation rate near the bottom of the radiative zone and the buoyancy radius for a large number of $\gamma$ Dor stars (see Van Reeth et al., 2015, 2016; Ouazzani et al., 2017; Van Reeth et al., 2018; Saio et al., 2018; Christophe et al., 2018; Mombarg et al., 2019; Ouazzani et al., 2019; Li et al., 2019, 2020; Takata et al., 2020), which led to unprecedented constraints on the models of angular momentum transport (Ouazzani et al., 2019; Li et al., 2020). These seismic analysis used an approximation of the rotation effects on gravity modes called the Traditional Approximation of Rotation (TAR, see Unno et al., 1989; Lee & Saio, 1989, 1997; Townsend, 2003; Bouabid et al., 2013). While generally convenient for modes confined in radiative zones (Ballot et al., 2012; Ouazzani et al., 2017, 2020), the TAR cannot describe modes oscillating in the convective cores of $\gamma$ Dor stars.
@@
Using an oscillation code which includes a full treatment of the Coriolis force, Ouazzani et al. (2020) showed the existence of gravito-inertial modes resonantly interacting with pure inertial modes confined in the convective core. In the otherwise smooth curve representing the period spacing between modes of consecutive order $\Delta P$ as a function of the period $P$, this interaction produces observable dips localised near the period of the inertial mode. It constitutes a major breakthrough since such a phenomenon can help probing the convective core of $\gamma$ Dor stars. Saio et al. (2021) analysed dips in 16 $\gamma$ Dor stars observed by Kepler and estimated their core rotation rate. Tokuno & Takata (2022) proposed a theoretical model of the shape of resonant dips in the $\Delta P$ $P$ curve. In this article, we study numerically the resonance between gravito-inertial and pure inertial modes of the convective core with the aim of extracting as much information as possible on the convective core from seismic data. For the different series of observed gravito-inertial modes (e.g., Li et al., 2020), we study the evolution of the dips with the rotation rate and the evolutionary stage of the star. This leads us to develop seismic diagnostic tools usable on observational data. We also test the model of Tokuno & Takata (2022) against our numerical results which leads us to construct a new model which better compares with these numerical results.
@@
In the following, we first present the stellar models and the oscillation code we used and discuss the modes identification in Sect. 2. We then present the results of our numerical computations for three series of modes at different rotation rates and evolutionary stages in Sect. 3. In Sect. 4, we provide an empirical model that relates the shape and position of dips to the inner structure of the stars. We discuss the model of Tokuno & Takata (2022) and present our new model in Sect. 5. We then move to the discussion and conclusion in Sect. 6.
@@ 2 Method
In this section we describe the oscillation code and the stellar models we used to study properties of the dip. We then explain how oscillation modes are identified and selected and finally we treat the question of numerical resolution.
@@ 2.1 Stellar models
In this subsection, we present the three main sequence star models we used for our study. They have been computed using the CESAM code (Morel, 1997; Morel & Lebreton, 2008). Their main characteristics are shown in Table 1. They are named after the CLES models used in Ouazzani et al. (2020) as they share similar characteristics. For these models, we adopted a solar metal mixture (Asplund et al., 2009), an initial helium fraction $Y=0.27$ and a initial metallicity $Z=0.02$. We used opacity tables from OPAL (Iglesias & Rogers, 1996), completed at low temperature with tables from Ferguson et al. (2005). We used the OPAL2005 equation of state (Rogers & Nayfonov, 2002) and the nuclear reaction rates from the NACRE collaboration (Angulo et al., 1999) except for the $\rm{}^{14}N(p,\gamma){}^{15}O$ reaction, for which we used the LUNA reaction rate given in Imbriani et al. (2004). Convection was treated using the mixing-length theory (Böhm-Vitense, 1958) with a mixing-length parameter $\alpha_{\mathrm{MLT}}=1.70$, close to a solar calibration. We included turbulent diffusion with a constant diffusion coefficient $D_{\mathrm{t}}=700\>\mathrm{cm^{2}\,s^{-1}}$. The physical prescriptions adopted for these models are thus very similar to the ones used in Ouazzani et al. (2020).
The three main-sequence models seek to cover the instability strip of $\gamma$ Dor stars (Bouabid et al., 2013). The model labelled ”1z” describes a young star of $1.40\>\mathrm{M_{\odot}}$ close to the Zero-Age Main Sequence (ZAMS) (age of $180$ My from the ZAMS, $X_{\rm c}=0.68$, $X_{\rm c}$ being the mass fraction of hydrogen in the core). The model labelled ”2m” describes an evolved mid-main sequence star of $1.60\>\mathrm{M_{\odot}}$ (age of $1800$ My from the ZAMS, $X_{\rm c}=0.38$). The model labelled ”3t” describes a star of $1.86\>\mathrm{M_{\odot}}$ at the end of the main sequence (age of $1480$ My from the ZAMS, $X_{\rm c}=0.05$).
Table 1: Characteristics of the 1z, 2m and 3t star models computed by CESAM.
@@
The Brunt-Väisälä frequency $N$ is an important parameter in our study since gravity modes depend directly on it. Moreover, the resonance with inertial modes in the convective core is greatly affected by the behaviour of $N$ at the bottom of the radiative zone, as we will demonstrate later. As shown in Fig. 1, the three models have different radial profiles of the Brunt-Väisälä frequency. They all present a discontinuity at the convective core/radiative zone interface, even the model that just leaves the ZAMS (model 1z). This is due to the chemical composition gradient that develops on the outer edge of the convective core as the star evolves on the main-sequence. This gradient, and thus the jump of the Brunt-Väisälä frequency, is sensitive to mixing processes such as rotational mixing, overshooting and microscopic diffusion. These mixing processes are modelled here through the turbulent diffusivity $D_{\rm t}$ mentioned above.
@@ 2.2 Oscillation code
To calculate oscillation modes and frequencies, we use the code TOP (Two-dimensional Oscillation Program, Reese, 2006; Reese et al., 2009), which we describe in this subsection.
We consider time-harmonic adiabatic small perturbations of a uniformly rotating spherical star. Thanks to the axial symmetry, the solutions are proportional to $e^{i(m\phi+\omega t)}$ where $m$ is the azimuthal number and $\omega$ is the mode frequency. The governing equations are:
$\rho+\@vec{\xi}\cdot\@vec{\nabla}\rho_{0}+\rho_{0}\@vec{\nabla}\cdot\@vec{\xi}=0,$ (1)
$2i\rho_{0}\@vec{\Omega}\times[\omega+m\Omega]\@vec{\xi}-[\omega+m\Omega]^{2}\rho_{0}\@vec{\xi}=-\@vec{\nabla}p+\frac{\@vec{\nabla}P_{0}}{\rho_{0}}\rho-\rho_{0}\@vec{\nabla}\psi,$ (2)
$p-c_{0}^{2}\rho+\@vec{\xi}\cdot\left(\@vec{\nabla}P_{0}-c_{0}^{2}\@vec{\nabla}\rho_{0}\right)=0,$ (3)
${\Delta}\psi-4\pi G\rho=0,$ (4)
where ${\xi}$, $p$, $\rho$ and $\psi$ are the amplitude of the Lagrangian displacement, the Eulerian perturbations of pressure, density and gravitational potential respectively, $\rho_{0}$, $P_{0}$ and $c_{0}$ are the density, the pressure and the sound velocity of the star model respectively with $c_{0}=\sqrt{\Gamma_{1}p_{0}/\rho_{0}}$, $\Gamma_{1}$ being the first adiabatic index. The star rotation rate is $\Omega$ and $G$ is the gravitational constant. The mode frequency in the inertial frame $\omega$ is related to the mode frequency in the co-rotating frame $\omega_{\mathrm{co}}$ through
@@
The governing equations are discretised using a fourth-order finite difference method in the radial direction (for details on the scheme, see Reese, 2013) and spectral decomposition on spherical harmonics in the horizontal directions. After discretisation, an algebraic eigenvalue problem is obtained and is solved using the Arnoldi-Chebyshev algorithm (Chatelin, 1988; Reese et al., 2009). Due the axial and equatorial symmetries of the problem, independent eigenvalue problems are solved for a given azimuthal order $m$ and a given parity with respect to the equator.
@@ 2.3 Mode identification
In practice, TOP finds a certain number of modes (typically 2 to 8) around a given frequency, for a prescribed resolution. These modes have the same equatorial parity and azimuthal order but modes associated to different degrees and radial orders are computed simultaneously. Unlike with 1D calculations, identification is thus non-trivial. Additionally, modes unresolved at the chosen resolution can be among the computed ones.
The series of gravito-inertial modes we want to compute are modes with the same degree $\ell$ and azimuthal number $m$ but different radial orders $n$. The degree $\ell$ corresponds to the associated Legendre polynomial $P_{\ell}^{m}$ that characterises the mode at zero rotation but remains relevant to label these same modes at higher rotation rates. For identification, we use the fact that the modes of the series are recognisable by their radial and latitudinal structure. We apply three criteria on the perturbed pressure $p$ to select the desired modes among the computed ones: two are related to the radial and latitudinal profiles while the third one takes care of miscalculated modes. Using the asymptotic Wentzel-Kramers-Brillouin (WKB) formulation of the TAR (see Unno et al., 1989), the number of radial nodes $n$ of a ($\ell,m$) gravito-inertial mode can be estimated by
@@
with $P_{\mathrm{co}}$ the period of the mode in the co-rotating frame, $\Pi_{0}$ the buoyancy radius, $\Lambda_{\ell}^{m}$ the eigenvalue associated with the Hough function $H_{\ell}^{m}$ describing the latitudinal part of the mode in the TAR approximation (see Appendix D.2) and $\epsilon_{g}$ being a small offset. We use this estimate of $n$ to remove modes with radial orders in the radiative zone noticeably higher than expected. Additionally, we compute the correlation of the mode latitudinal profiles in the radiative zone with the corresponding Hough function and reject modes with weak correlation coefficients. Finally, we exclude badly resolved and spurious modes. This includes modes with an extremely high amplitude in the core or at the surface as well as modes showing strong variations between consecutive radial grid points.

### v2

# Properties of observable mixed inertial and gravito-inertial modes in $\gamma$ Doradus stars

@@ Abstract
Context. The space missions Kepler and TESS provided a large number of highly detailed time series for main-sequence stars, including $\gamma$ Doradus stars. Additionally, numerous $\gamma$ Doradus stars are to be observed in the near future thanks to the upcoming PLATO mission. In $\gamma$ Doradus stars, gravito-inertial modes in the radiative zone and inertial modes in the convective core can interact resonantly, which translates into the appearance of dip structures in the period spacing of modes. Those dips are information-rich, as they are related to the star core characteristics.
Aims. Our aim is to characterise these dips according to stellar properties and thus to develop new seismic diagnostic tools to constrain the internal structure of $\gamma$ Doradus stars, especially their cores.
Methods. We used the two-dimensional oscillation code TOP to compute sectoral prograde and axisymmetric dipolar modes in $\gamma$ Doradus stars at different rotation rates and evolutionary stages. We then characterised the dips we obtained by their width and location on the period spacing diagram.
Results. We found that the width and the location of the dips depend quasi-linearly on the ratio of the rotation rate and the Brunt-Väisälä frequency at the core interface. This allowed us to determine empirical relations between the width and location of dips as well as the resonant inertial mode frequency in the core and the Brunt-Väisälä frequency at the interface between the convective core and the radiative zone. We propose an approximate theoretical model to support and discuss these empirical relations.
Conclusions. The empirical relations we established could be applied to dips observed in data, which would allow for the estimation of frequencies of resonant inertial modes in the core and of the Brunt-Väisälä jump at the interface between the core and the radiative zone. As those two parameters are both related to the evolutionary stage of the star, their determination could lead to more accurate estimations of stellar ages.
@@ 1 Introduction
Asteroseismology has proven to be a crucial tool in unravelling the mysteries of stellar evolution and determining the ages and evolutionary stages of stars (e.g. Lebreton et al., 2014; Aerts, 2021, and references therein). Estimating the evolutionary stages of stars is particularly significant to addressing the complex question of angular momentum transport within stars (Aerts et al., 2019).
@@
Among pulsating stars, $\gamma$ Doradus (hereafter $\gamma$ Dor) stars are main-sequence stars that typically have a mass of 1.4–2.0 M and a radius of 1.4–2.3 R (Kaye et al., 1999; Uytterhoeven et al., 2011). Their structure is characterised by a convective core, a large radiative zone, and a small sub-surface convective envelope. They exhibit non-radial gravity modes oscillating with a period of 0.3 to 3 d (Kaye et al., 1999). Thanks to high-precision photometry provided by space-borne instruments, especially by the Kepler mission (Borucki et al., 2010) and the Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2014), a large quantity of good quality data is now available, and it has enabled the study of numerous $\gamma$ Dor stars.
@@
Gravito-inertial modes have been identified in 611 $\gamma$ Dor stars observed by Kepler (Li et al., 2020) and in 106 $\gamma$ Dor stars observed by TESS (Garcia et al., 2022). These observed modes have enabled both the rotation rate near the bottom of the radiative zone and the buoyancy radius for a large number of $\gamma$ Dor stars to be determined (see Van Reeth et al., 2015, 2016; Ouazzani et al., 2017; Van Reeth et al., 2018; Saio et al., 2018; Christophe et al., 2018; Mombarg et al., 2019; Ouazzani et al., 2019; Li et al., 2019, 2020; Takata et al., 2020), which has led to unprecedented constraints on the models of angular momentum transport (Ouazzani et al., 2019; Li et al., 2020). These seismic analyses used an approximation of the rotation effects on gravity modes called the traditional approximation of rotation (TAR; see Unno et al., 1989; Lee & Saio, 1989, 1997; Townsend, 2003; Bouabid et al., 2013). While generally convenient for modes confined in radiative zones (Ballot et al., 2012; Ouazzani et al., 2017, 2020), the TAR cannot describe modes oscillating in the convective cores of $\gamma$ Dor stars.
@@
Using an oscillation code that includes a full treatment of the Coriolis force, Ouazzani et al. (2020) showed the existence of gravito-inertial modes resonantly interacting with pure inertial modes confined in the convective core. In the otherwise smooth curve representing the period spacing between modes of consecutive order $\Delta P$ as a function of the period $P$, this interaction produces observable dips localised near the period of the inertial mode. This discovery constitutes a major breakthrough since such a phenomenon can help probe the convective core of $\gamma$ Dor stars. Saio et al. (2021) analysed dips in 16 $\gamma$ Dor stars observed by Kepler and estimated their core rotation rate. Moreover, Tokuno & Takata (2022) proposed a theoretical model of the shape of resonant dips in the $\Delta P$ $P$ curve.
In this article, we numerically study the resonance between gravito-inertial and pure inertial modes of the convective core with the aim of extracting as much information as possible on the convective core from seismic data. For the different series of observed gravito-inertial modes (e.g. Li et al., 2020), we studied the evolution of the dips with the rotation rate and the evolutionary stage of the star. This led us to develop seismic diagnostic tools that are usable on observational data. We also tested the model of Tokuno & Takata (2022) against our numerical results, which led us to construct a new model that better compares with these numerical results.
@@
In the following, we first present the stellar models and the oscillation code we used and discuss the mode identification in Sect. 2. We then present the results of our numerical computations for three series of modes at different rotation rates and evolutionary stages in Sect. 3. In Sect. 4, we provide an empirical model that relates the shape and position of dips to the inner structure of the stars. We discuss the model of Tokuno & Takata (2022) and present our new model in Sect. 5. We then move to a discussion and the conclusion of paper in Sect. 6.
@@ 2 Method
In this section we describe the oscillation code and the stellar models we used to study properties of the dip. We then explain how oscillation modes are identified and selected, and finally we treat the question of numerical resolution.
@@ 2.1 Stellar models
In this subsection, we present the three main-sequence star models we used for our study. They were computed using the CESAM code (Morel, 1997; Morel & Lebreton, 2008). Their main characteristics are shown in Table 1. They are named after the CLES models used in Ouazzani et al. (2020), as they share similar characteristics. For these models, we adopted a solar metal mixture (Asplund et al., 2009), an initial helium fraction $Y=0.27$, and a initial metallicity $Z=0.02$. We used opacity tables from OPAL (Iglesias & Rogers, 1996) completed at low temperature with tables from Ferguson et al. (2005). We used the OPAL2005 equation of state (Rogers & Nayfonov, 2002) and the nuclear reaction rates from the NACRE collaboration (Angulo et al., 1999) except for the $\rm{}^{14}N(p,\gamma){}^{15}O$ reaction, for which we used the LUNA reaction rate given in Imbriani et al. (2004). Convection was treated using the mixing-length theory (Böhm-Vitense, 1958) with a mixing-length parameter $\alpha_{\mathrm{MLT}}=1.70$, close to a solar calibration. We included turbulent diffusion with a constant diffusion coefficient of $D_{\mathrm{t}}=700\>\mathrm{cm^{2}\,s^{-1}}$. The physical prescriptions adopted for these models are thus very similar to the ones used in Ouazzani et al. (2020).
The three main-sequence models seek to cover the instability strip of $\gamma$ Dor stars (Bouabid et al., 2013). The model labelled ‘1z’ describes a young star of $1.40\>\mathrm{M_{\odot}}$ close to the zero age main sequence (ZAMS; age of $180$ My from the ZAMS, $X_{\rm c}=0.68$, with $X_{\rm c}$ being the mass fraction of hydrogen in the core). The model labelled ‘2m’ describes an evolved mid-main-sequence star of $1.60\>\mathrm{M_{\odot}}$ (age of $1800$ My from the ZAMS, $X_{\rm c}=0.38$). The model labelled ‘3t’ describes a star of $1.86\>\mathrm{M_{\odot}}$ at the end of the main sequence (age of $1480$ My from the ZAMS, $X_{\rm c}=0.05$).
Table 1: Characteristics of the 1z, 2m, and 3t star models computed by CESAM.
@@
The Brunt-Väisälä frequency $N$ is an important parameter in our study since gravity modes depend directly on it. Moreover, the resonance with inertial modes in the convective core is greatly affected by the behaviour of $N$ at the bottom of the radiative zone, as we demonstrate later. As shown in Fig. 1, the three models have different radial profiles of the Brunt-Väisälä frequency. They all present a discontinuity at the convective core-radiative zone interface, even the model that has just left the ZAMS (model 1z). This is due to the chemical composition gradient that develops on the outer edge of the convective core as the star evolves on the main sequence. This gradient, and thus the jump of the Brunt-Väisälä frequency, is sensitive to mixing processes such as rotational mixing, overshooting and microscopic diffusion. These mixing processes are modelled here through the turbulent diffusivity $D_{\rm t}$ mentioned above.
@@ 2.2 Oscillation code
To calculate the oscillation modes and frequencies, we used the code TOP (Two-dimensional Oscillation Program, Reese, 2006; Reese et al., 2009), which we describe in this subsection.
We considered time-harmonic adiabatic small perturbations of a uniformly rotating spherical star. Thanks to the axial symmetry, the solutions are proportional to $e^{i(m\phi+\omega t)}$, where $m$ is the azimuthal number and $\omega$ is the mode frequency. The governing equations are
$\rho+\vec{\xi}\cdot\vec{\nabla}\rho_{0}+\rho_{0}\vec{\nabla}\cdot\vec{\xi}=0,$ (1)
$2i\rho_{0}\vec{\Omega}\times[\omega+m\Omega]\vec{\xi}-[\omega+m\Omega]^{2}\rho_{0}\vec{\xi}=-\vec{\nabla}p+\frac{\vec{\nabla}P_{0}}{\rho_{0}}\rho-\rho_{0}\vec{\nabla}\psi,$ (2)
$p-c_{0}^{2}\rho+\vec{\xi}\cdot\left(\vec{\nabla}P_{0}-c_{0}^{2}\vec{\nabla}\rho_{0}\right)=0,$ (3)
$\vec{\Delta}\psi-4\pi G\rho=0,$ (4)
where $\vec{\xi}$, $p$, $\rho$, and $\psi$ are the amplitude of the Lagrangian displacement, the Eulerian perturbations of pressure, density, and gravitational potential, respectively, and where $\rho_{0}$, $P_{0}$, and $c_{0}$ are the density, the pressure, and the sound velocity of the star model, respectively, with $c_{0}=\sqrt{\Gamma_{1}p_{0}/\rho_{0}}$, $\Gamma_{1}$ being the first adiabatic index. The star rotation rate is $\Omega$, and $G$ is the gravitational constant. The mode frequency in the inertial frame $\omega$ is related to the mode frequency in the co-rotating frame $\omega_{\mathrm{co}}$ through
@@
The governing equations were discretised using a fourth-order finite difference method in the radial direction (for details on the scheme, see Reese, 2013) and spectral decomposition on spherical harmonics in the horizontal directions. After discretisation, an algebraic eigenvalue problem was obtained and solved using the Arnoldi-Chebyshev algorithm (Chatelin, 1988; Reese et al., 2009). Due the axial and equatorial symmetries of the problem, independent eigenvalue problems were solved for a given azimuthal order $m$ and a given parity with respect to the equator.
@@ 2.3 Mode identification
In practice, TOP finds a certain number of modes (typically two to eight) around a given frequency for a prescribed resolution. These modes have the same equatorial parity and azimuthal order, but modes associated with different degrees and radial orders are computed simultaneously. Unlike with 1D calculations, identification is thus non-trivial. Additionally, modes unresolved at the chosen resolution can be among the computed ones.
The series of gravito-inertial modes we computed for this work are the modes with the same degree $\ell$ and azimuthal number $m$ but different radial orders $n$. The degree $\ell$ corresponds to the associated Legendre polynomial $P_{\ell}^{m}$ that characterises the mode at zero rotation but remains relevant to label these same modes at higher rotation rates. For identification, we used the fact that the modes of the series are recognisable by their radial and latitudinal structure. We applied three criteria on the perturbed pressure $p$ to select the desired modes among the computed ones; two are related to the radial and latitudinal profiles, while the third criterion takes care of miscalculated modes. Using the asymptotic Wentzel-Kramers-Brillouin (WKB) formulation of the TAR (see Unno et al., 1989), the number of radial nodes $n$ of a ($\ell,m$) gravito-inertial mode can be estimated by
@@
with $P_{\mathrm{co}}$ as the period of the mode in the co-rotating frame, $\Pi_{0}$ as the buoyancy radius, $\Lambda_{\ell}^{m}$ as the eigenvalue associated with the Hough function $H_{\ell}^{m}$ describing the latitudinal part of the mode in the TAR (see Appendix D.2), and $\epsilon_{g}$ as a small offset. We used this estimate of $n$ to remove modes with radial orders in the radiative zone noticeably higher than expected. Additionally, we computed the correlation of the mode latitudinal profiles in the radiative zone with the corresponding Hough function and rejected modes with weak correlation coefficients. Finally, we excluded badly resolved and spurious modes. This included modes with an extremely high amplitude in the core or at the surface as well as modes showing strong variations between consecutive radial grid points.